# Efficient quantum Gibbs samplers with Kubo–Martin–Schwinger detailed balance condition

### v2

@@ 4. Recovery of the Gibbs sampler in [CKG23]
We have seen from (2.16) that Davies generator (2.17) without the coherent term (i.e., Lindbladian with $\sigma_{\beta}$-GNS DBC) corresponds to the algorithmic Lindbladian (1.2) with $\widehat{f}(\omega)=\delta(\omega)/2\pi$ and $G=0$. As emphasized in 7, such a choice of Dirac delta filtering function for the frequency makes it hard to approximate GNS detailed balanced Lindblad dynamic. Chen et al. [CKBG23, Theorem I.3] introduced a Gaussian smoothed version by taking $f$ as
@@
(4.2) $\mathcal{L}^{{\dagger}}[\rho]$ $=-i[G,\rho]+\sum_{a\in\mathcal{A}}\int^{\infty}_{-\infty}\gamma(\omega)\left(\widehat{A}_{f}^{a}(\omega)\rho\left(\widehat{A}_{f}^{a}(\omega)\right)^{\dagger}-\frac{1}{2}\left\{\left(\widehat{A}_{f}^{a}(\omega)\right)^{\dagger}\widehat{A}_{f}^{a}(\omega),\rho\right\}\right)\,\mathrm{d}\omega$
@@
$|\alpha_{\nu,\nu^{\prime}}|\leq C\left\lVert\gamma\right\rVert_{L^{\infty}(\mathbb{R})}\beta\int^{\infty}_{-\infty}e^{-\frac{\beta^{2}((\omega-\nu)^{2}+(\omega-\nu^{\prime})^{2})}{4}}\mathrm{d}\omega\leq C\left\lVert\gamma\right\rVert_{L^{\infty}(\mathbb{R})}e^{-\frac{\beta^{2}(\nu-\nu^{\prime})}{8}}\,,$
@@
which is always supported on $[-S,S]^{2}$ independent of $\beta$ by 15. We refer the readers to Fig. 3 below for an illustration of the pattern of $\alpha_{\nu,\nu^{\prime}}$ for various Gibbs samplers.
@@
Recall 20 and note that in [CKG23], both choices of $\gamma^{(g)}$ and $\gamma^{(m)}$ in Eqs. 4.3 and 4.4 give the linear dependence of the total Hamiltonian simulation time on $\beta$. The discrepancy in complexity between our approach and theirs arises from the difference in the construction of jump operators and coherent terms. The discussion above shows that when $|\nu-\nu^{\prime}|=\Omega(1/\beta)$, $\alpha_{\nu,\nu^{\prime}}\ll 1$ (see Eq. 4.13). Consequently, in (4.2), the energy transition $\nu$ in $A_{\nu}$ is always $\mathcal{O}(1/\beta)$ close to the energy transition $\nu^{\prime}$ in $(A^{a}_{\nu^{\prime}})^{\dagger}$. This property ensures that the norm of the coherent term $G$ in [CKG23] and the normalization block encoding constant does not increase linearly in $\beta$. This differs from our case, where $\{\alpha_{\nu,\nu^{\prime}}\}$ always includes a $\Omega(1)$-sized principal submatrix and the dynamics allows different energy transition terms $A_{\nu}\rho(A^{a}_{\nu^{\prime}})^{\dagger}$ even when $|\nu-\nu^{\prime}|\gg\Omega(1/\beta)$. Therefore, our coherent $G$ has much more cross terms in the expansion, and the normalization constant for the block encoding increases linearly in $\beta$ (3.31). This introduces an additional $\beta$ factor in our complexity result (see 19).
@@ Appendix C Gevrey functions
$B_{k,j}=\sum_{\begin{subarray}{c}1\leq i\leq k\,,\,q_{i}\in\mathbb{N}\\ \sum^{k}_{i=1}iq_{i}=k\\ \sum^{k}_{i=1}q_{i}=j\end{subarray}}\frac{k!}{q_{1}!q_{2}!\cdots q_{k-j+1}!}\prod^{k-j+1}_{i=1}\left(\frac{x_{i}}{i!}\right)^{q_{i}}\,.$
@@
By 14, we have, for every $d$-tuple of nonnegative integers $\alpha$ with $|\alpha|=\sum_{i}\left\lvert\alpha_{i}\right\rvert$,
@@
Appendix D Quadrature error analysis
In this section, for notational simplicity, sometimes we absorb the generic constant $C_{s}$ depending on the parameter $s$ of the filtering function $q(\nu)$ in Eq. 3.12 into $\mathcal{O}$. We first study the decaying and integrable property of the functions $f$ and $g(t,t^{\prime})$.
@@ Appendix D Quadrature error analysis
Let $f(t)$ and $g(t,t^{\prime})$ be the functions defined in (3.3) and (3.6) with a weighting function $q(\nu)$ satisfying 15. Then, it holds that
(D.1) $|f(t)|\leq\frac{A^{2}_{q}S}{\pi}\exp\left(\frac{es}{2}-\frac{s}{(\beta A_{u}+A_{w})^{1/s}e}|t|^{1/s}\right)\,,$
@@
(D.2) $\left|g(t,t^{\prime})\right|\leq\frac{A^{4}_{q}S^{2}}{2\pi^{2}}\exp\left(es-\frac{s}{(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}\left(\sqrt{t^{2}+(t^{\prime})^{2}}\right)^{1/s}\right)\,.$
@@
(D.3) $\int_{\mathbb{R}}|f(t)|\,\mathrm{d}t=\mathcal{O}\left((A_{q}C_{1,u}+A^{2}_{q}SA_{w})\log(\beta A_{u}+A_{w})\right)\,,$
@@
(D.4) $\iint_{\mathbb{R}^{2}}|g(t,t^{\prime})|\,\mathrm{d}t\,\mathrm{d}t^{\prime}=\mathcal{O}\left((A_{q}^{2}C_{1,u}^{2}+A_{q}^{4}S^{2}A_{w}^{2}+A_{q}^{3}C_{1,u}SA_{w}+A_{q}^{4}\beta S)\log^{2}(\beta A_{u}+A_{w}+\beta)\right)\,.$
@@
We recall the definitions of $f(t)$ and $g(t,t^{\prime})$:
(D.5) $f(t)=\frac{1}{2\pi}\int_{\mathbb{R}}u(\beta\nu)w(\nu)e^{-\beta\nu/4}e^{-it\nu}\,\mathrm{d}\nu\,,$
@@
(D.6) $g(t,t^{\prime})=\frac{1}{(2\pi)^{2}}\iint_{\mathbb{R}^{2}}\frac{\tanh(\beta(\nu^{\prime}-\nu)/4)e^{-\beta(\nu+\nu^{\prime})/4}}{2i}u(\beta\nu)w(\nu)\overline{u(\beta\nu^{\prime})w(\nu^{\prime})}e^{-i(\nu t-\nu^{\prime}t^{\prime})}\,\mathrm{d}\nu\,\mathrm{d}\nu^{\prime}\,.$
Exponential decay of $f$. Thanks to 26 with 15, we have
$\widehat{f}(\nu)=u(\beta\nu)w(\nu)e^{-\beta\nu/4}\in\mathcal{G}^{s}_{A^{2}_{q},\beta A_{u}+A_{w}}(\mathbb{R})\quad\text{with}\quad\mathrm{supp}\big{(}\widehat{f}\big{)}\in[-S,S]\,.$
@@
(D.7) $\left|f(t)\right|\leq\frac{A^{2}_{q}S}{\pi}\exp\left(\frac{es}{2}-\frac{s|t|^{1/s}}{(\beta A_{u}+A_{w})^{1/s}e}\right)\,.$
@@
(D.8) $\tanh((\nu^{\prime}-\nu)/4)\in\mathcal{G}^{1}_{1,1}(\mathbb{R}^{2})\,.$
It follows from 26 with
$(2i)\widehat{g}(\nu,\nu^{\prime})=\tanh(\beta(\nu^{\prime}-\nu)/4)e^{-\beta(\nu+\nu^{\prime})/4}u(\beta\nu)w(\nu)\overline{u(\beta\nu^{\prime})w(\nu^{\prime})}$
that
$(2i)\widehat{g}(\nu,\nu^{\prime})\in\mathcal{G}^{s}_{A^{4}_{q},2\beta A_{u}+2A_{w}+\beta}\quad\text{with}\quad\mathrm{supp}\big{(}\widehat{g}\big{)}\in[-S,S]^{2}\,.$
Again, applying 29 with $d=2$ and $|\Omega|=4S^{2}$ gives
(D.9) $\left|g(t,t^{\prime})\right|\leq\frac{A^{4}_{q}S^{2}}{2\pi^{2}}\exp\left(es-\frac{s}{(2\beta A_{u}+2A_{w}+\beta)^{1/s}e}\left(\sqrt{t^{2}+(t^{\prime})^{2}}\right)^{1/s}\right)\,,$
@@
Estimate the integral of $f$. We note from 15 that the derivative of $u(\beta\nu)w(\nu)e^{-\beta\nu/4}$ is $L^{1}$-integrable. It follows that
$itf(t)=\frac{1}{2\pi}\int_{\mathbb{R}}\frac{\mathrm{d}}{\mathrm{d}\nu}\left(u(\beta\nu)w(\nu)e^{-\beta\nu/4}\right)e^{-it\nu}\,\mathrm{d}\nu\,,$
@@
$2\pi\left\lVert tf(t)\right\rVert_{L^{\infty}(\mathbb{R})}$ $\leq\left\lVert(u(\beta\nu)e^{-\beta\nu/4})^{(1)}\right\rVert_{L^{1}(\mathbb{R})}\left\lVert w\right\rVert_{L^{\infty}(\mathbb{R})}+\left\lVert u(\beta\nu)e^{-\beta\nu/4}\right\rVert_{L^{\infty}(\mathbb{R})}\left\lVert w^{(1)}\right\rVert_{L^{1}(\mathbb{R})}$
$\leq A_{q}C_{1,u}+2SA^{2}_{q}A_{w}\,,$
by the invariance of $\left\lVert(u(\beta\nu)e^{-\beta\nu/4})^{(1)}\right\rVert_{L^{1}(\mathbb{R})}$ in $\beta$. Here, we use $w\in\mathcal{G}^{s}_{A_{q},A_{\omega}}$ and $\mathrm{supp}(w)=[-S,S]$ to obtain $\left\lVert w^{(1)}\right\rVert_{L^{1}(\mathbb{R})}\leq 2SA_{q}A_{\omega}$. In addition, because $\left\lVert u(\beta\nu)e^{-\beta\nu/4}\right\rVert_{L^{\infty}(\mathbb{R})}\leq A_{q}$, there holds
$\left\lVert f\right\rVert_{L^{\infty}(\mathbb{R})}\leq\frac{1}{2\pi}\int_{[-S,S]}\left|u(\beta\nu)w(\nu)e^{-\beta\nu/4}\right|\,\mathrm{d}\nu\leq\frac{1}{\pi}A_{q}^{2}S\,.$
Thus, we readily have, for any $T>1$,
(D.10) $\int_{-T}^{T}|f(t)|\,\mathrm{d}t$ $\leq\frac{2}{\pi}A_{q}^{2}S+2\int_{1}^{T}|f(t)|\,\mathrm{d}t$
$\leq\frac{2}{\pi}A_{q}^{2}S+2\frac{A_{q}C_{1,u}+2SA^{2}_{q}A_{w}}{2\pi}\log T=\mathcal{O}\left((A_{q}C_{1,u}+SA^{2}_{q}A_{w})\log T\right)\,.$
@@
(D.11) $\int_{T}^{\infty}\left|f(t)\right|\,\mathrm{d}t=\mathcal{O}\left(A^{2}_{q}S\int_{T^{1/s}}^{\infty}u^{s-1}\exp\left(-\frac{s}{(\beta A_{u}+A_{w})^{1/s}e}u\right)\,\mathrm{d}u\right)\,,\quad T>0\,.$
@@
(D.12) $T_{f}(A_{u},A_{w}):=\inf\left\{T>0\,;\ u^{s-1}\leq\exp\left(\frac{s}{2(\beta A_{u}+A_{w})^{1/s}e}u\right)\ \text{for any $u\geq T^{1/s}$}\right\},$
which satisfies the following asymptotics: as $\beta A_{u}+A_{w}\to\infty$,
$T_{f}(A_{u},A_{w})/\log(T_{f}(A_{u},A_{w}))=\Theta\left(\beta A_{u}+A_{w}\right)\,.$
@@
(D.13) $T_{f}(A_{u},A_{w})=\Theta\left((\beta A_{u}+A_{w})\log\left(\beta A_{u}+A_{w}\right)\right)\,.$
We can then estimate the integral by Eqs. D.10 and D.11, as well as Eq. D.13,
@@
(D.14) $\int_{\mathbb{R}}\left|f(t)\right|\,\mathrm{d}t$ $=\int_{-T^{k}_{f}}^{T^{k}_{f}}|f(t)|\,\mathrm{d}t+\mathcal{O}\left(A^{2}_{q}S\int_{T_{f}^{k/s}}^{\infty}\exp\left(-\frac{s}{2(\beta A_{u}+A_{w})^{1/s}e}u\right)\,\mathrm{d}u\right)$
$=\mathcal{O}\left((A_{q}C_{1,u}+A^{2}_{q}SA_{w})\log T_{f}+A^{2}_{q}S\underbrace{(\beta A_{u}+A_{w})^{1/s}\exp\left(-\frac{s}{2(\beta A_{u}+A_{w})^{1/s}e}T_{f}^{k/s}\right)}_{\text{$o(1)$ as $\beta A_{u}+A_{w}\to\infty$}}\right)$
$=\mathcal{O}\left((A_{q}C_{1,u}+A^{2}_{q}SA_{w})\log(\beta A_{u}+A_{w})\right),$
for any positive $k>1$. This gives Eq. D.3.

### v4

@@ 4. Recovery of the Gibbs sampler in [CKG23]
[…]
@@
(4.2) $\mathcal{L}^{{\dagger}}(\rho)$ $=-i[G,\rho]+\sum_{a\in\mathcal{A}}\int^{\infty}_{-\infty}\gamma(\omega)\left(\widehat{A}_{f}^{a}(\omega)\rho\left(\widehat{A}_{f}^{a}(\omega)\right)^{\dagger}-\frac{1}{2}\left\{\left(\widehat{A}_{f}^{a}(\omega)\right)^{\dagger}\widehat{A}_{f}^{a}(\omega),\rho\right\}\right)\,\mathrm{d}\omega$
@@
$|\alpha_{\nu,\nu^{\prime}}|\leq C\left\lVert\gamma\right\rVert_{L^{\infty}(\mathbb{R})}\beta\int^{\infty}_{-\infty}e^{-\frac{\beta^{2}((\omega-\nu)^{2}+(\omega-\nu^{\prime})^{2})}{4}}\mathrm{d}\omega\leq C\left\lVert\gamma\right\rVert_{L^{\infty}(\mathbb{R})}e^{-\frac{\beta^{2}(\nu-\nu^{\prime})^{2}}{8}}\,,$
@@
which is always supported on $[-S,S]^{2}$ independent of $\beta$ by 13. We refer the readers to Fig. 3 below for an illustration of the pattern of $\alpha_{\nu,\nu^{\prime}}$ for various Gibbs samplers.
@@
The discussion above shows that when $|\nu-\nu^{\prime}|=\Omega(1/\beta)$, $\alpha_{\nu,\nu^{\prime}}\ll 1$ (see Eq. 4.13). Consequently, in (4.2), the energy transition $\nu$ in $A_{\nu}$ is always $\mathcal{O}(1/\beta)$ close to the energy transition $\nu^{\prime}$ in $(A^{a}_{\nu^{\prime}})^{\dagger}$. This differs from our case, where $\{\alpha_{\nu,\nu^{\prime}}\}$ always includes a $\Omega(1)$-sized principal submatrix and the dynamics allows different energy transition terms $A_{\nu}\rho(A^{a}_{\nu^{\prime}})^{\dagger}$ even when $|\nu-\nu^{\prime}|\gg\Omega(1/\beta)$.
@@ Appendix C Gevrey functions
$B_{k,j}(z_{1},z_{2},\dots,z_{k-j+1})=\sum_{\begin{subarray}{c}1\leq i\leq k\,,\,q_{i}\in\mathbb{N}\\ \sum^{k}_{i=1}iq_{i}=k\\ \sum^{k}_{i=1}q_{i}=j\end{subarray}}\frac{k!}{q_{1}!q_{2}!\cdots q_{k-j+1}!}\prod^{k-j+1}_{i=1}\left(\frac{z_{i}}{i!}\right)^{q_{i}}\,.$
@@
By 12, we have, for every $d$-tuple of nonnegative integers $\alpha$ with $|\alpha|=\sum_{i}\left\lvert\alpha_{i}\right\rvert$,
@@
Appendix D Error analysis and complexity
In this section, for notational simplicity, we often absorb the generic constant $C_{s}$ depending on the parameter $s$ of the weighing function $q(\nu)$ in Eq. 3.16 into $\mathcal{O}$. We first study the decaying and integrable properties of functions $f$ and $g$ involved in the jumps and the coherent term, respectively.
@@ Appendix D Error analysis and complexity
Let $f(t)$ and $g(t)$ be the functions defined in (3.4) and (3.7)-(3.9) with $q(\nu)$ and $\kappa(\nu)$ satisfying 13 and Eq. 3.17. Then, it holds that
(D.1) $|f(t)|\leq\frac{\xi^{2}_{q}S}{\pi}\exp\left(\frac{es}{2}-\frac{s}{e}\left(\frac{|t|}{\beta\xi_{u}+\xi_{w}/S}\right)^{1/s}\right)\,,$
@@
(D.2) $\left|g(t)\right|\leq\frac{2\xi_{q}S}{\pi}\exp\left(\frac{es}{2}-\frac{s}{e}\left(\frac{|t|}{\beta+\xi_{w}/4S}\right)^{1/s}\right)\,.$
@@
(D.3) $\int_{\mathbb{R}}|f(t)|\,\mathrm{d}t=\mathcal{O}\left((\xi_{q}C_{1,u}+\xi^{2}_{q}\xi_{w})\log((\beta\xi_{u}+\xi_{w}/S)\max\{S,1\})\right)\,,$
@@
(D.4) $\int_{\mathbb{R}}|g(t)|\,\mathrm{d}t=\mathcal{O}\left(\xi_{q}(1+\log((\beta+\xi_{w}/S)\max\{S,1\}))\right)\,.$
@@
We recall the definitions of $f(t)$ and $g(t)$:
(D.5) $f(t)=\frac{1}{2\pi}\int_{\mathbb{R}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u(\beta\nu)w(\nu/S)e^{-\beta\nu/4}}e^{-it\nu}\,\mathrm{d}\nu\,,$
@@
(D.6) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}g(t)=\frac{1}{2\pi}\int_{\mathbb{R}}-\frac{i}{2}\tanh\left(-\beta\nu/4\right)w(\nu/4S)e^{-i\nu t}\,\mathrm{d}\nu\,\mathrm{d}\nu^{\prime}\,.}$
Exponential decay of $f$. Thanks to 26 with 13, we have
$\widehat{f}(\nu)=u(\beta\nu)w(\nu/S)e^{-\beta\nu/4}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{G}^{s}_{\xi^{2}_{q},\beta\xi_{u}+\xi_{w}/S}(\mathbb{R})}\quad\text{with}\quad\mathrm{supp}\big{(}\widehat{f}\big{)}\in[-S,S]\,.$
@@
(D.7) $\left|f(t)\right|\leq\frac{\xi^{2}_{q}S}{\pi}\exp\left(\frac{es}{2}-\frac{s}{e}\left(\frac{|t|}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta\xi_{u}+\xi_{w}/S}}\right)^{1/s}\right)\,.$
@@
(D.8) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tanh\left(-\nu\beta/4\right)\in\mathcal{G}^{1}_{1,\beta}(\mathbb{R})\,.}$
It follows from 26 that
(D.9) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(2i)\widehat{g}(\nu):=\tanh\left(-\nu\beta/4\right)w\left(\nu/4S\right)\in\mathcal{G}^{s}_{\xi_{q},\beta+\xi_{w}/4S}\left(\mathbb{R}\right),\quad\mathrm{supp}(\widehat{g})=[-4S,4S]\,.}$
Again, applying 29 with $d=1$ and $|\Omega|=8S$ gives
(D.10) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left|g(t)\right|\leq\frac{2\xi_{q}S}{\pi}\exp\left(\frac{es}{2}-\frac{s}{e}\left(\frac{|t|}{\beta+\xi_{w}/4S}\right)^{1/s}\right)\,.}$
@@
Estimate the integral of $f$. We note from 13 that the derivative of $u(\beta\nu)w(\nu/S)e^{-\beta\nu/4}$ is $L^{1}$-integrable. It follows that
$itf(t)=\frac{1}{2\pi}\int_{\mathbb{R}}\frac{\mathrm{d}}{\mathrm{d}\nu}\left(u(\beta\nu)w(\nu/S)e^{-\beta\nu/4}\right)e^{-it\nu}\,\mathrm{d}\nu\,,$
@@
(D.11) $2\pi\left\lVert tf(t)\right\rVert_{L^{\infty}(\mathbb{R})}$ $\leq\|(u(\beta\nu)e^{-\beta\nu/4})^{(1)}\|_{L^{1}(\mathbb{R})}\|w(\nu/S)\|_{L^{\infty}(\mathbb{R})}$
$\qquad\qquad\qquad+\|u(\beta\nu)e^{-\beta\nu/4}|_{L^{\infty}(\mathbb{R})}\|(w(\nu/S))^{(1)}\|_{L^{1}(\mathbb{R})}$
$\leq\xi_{q}C_{1,u}+2\xi^{2}_{q}\xi_{w}\,,$
by the invariance of $\|(u(\beta\nu)e^{-\beta\nu/4})^{(1)}\|_{L^{1}(\mathbb{R})}$ in $\beta$. Here, we use $w\in\mathcal{G}^{s}_{\xi_{q},\xi_{w}}$ and $\mathrm{supp}(w)=[-1,1]$ to obtain $\|(w(\nu/S))^{(1)}\|_{L^{1}(\mathbb{R})}\leq 2\xi_{q}\xi_{w}$. In addition, because $\|u(\beta\nu)e^{-\beta\nu/4}\|_{L^{\infty}(\mathbb{R})}\leq\xi_{q}$, there holds
$\left\lVert f\right\rVert_{L^{\infty}(\mathbb{R})}\leq\frac{1}{2\pi}\int_{-S}^{S}\left|u(\beta\nu)w(\nu/S)e^{-\beta\nu/4}\right|\,\mathrm{d}\nu\leq\frac{1}{\pi}\xi_{q}^{2}S\,.$
Thus, we readily have, for any $T>\min\{1,1/S\}$,
(D.12) $\int_{-T}^{T}|f(t)|\,\mathrm{d}t$ $\leq\int_{|t|\leq\min\{1,1/S\}}|f(t)|\,\mathrm{d}t+\int_{\min\{1,1/S\}\leq|t|\leq T}|f(t)|\,\mathrm{d}t$
$\leq\frac{2}{\pi}\xi_{q}^{2}\min\{S,1\}+2\frac{\xi_{q}C_{1,u}+2\xi^{2}_{q}\xi_{w}}{2\pi}\log(T\max\{S,1\})$
$=\mathcal{O}\left((\xi_{q}C_{1,u}+\xi^{2}_{q}\xi_{w})\log(T\max\{S,1\})\right).$
@@
(D.13) $\int_{T}^{\infty}\left|f(t)\right|\,\mathrm{d}t=\mathcal{O}\left(\xi^{2}_{q}S\int_{T^{1/s}}^{\infty}u^{s-1}\exp\left(-\frac{s}{e({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\beta\xi_{u}+\xi_{w}/S)^{1/s}}}u\right)\,\mathrm{d}u\right)\,,\quad T>0\,.$
@@
(D.14) $T_{f}:=\inf\left\{T>0\,;\ u^{s-1}\leq\exp\left(\frac{s}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2e(\beta\xi_{u}+\xi_{w}/S)^{1/s}}}u\right)\ \text{for any $u\geq T^{1/s}$}\right\},$
which satisfies the following asymptotics:
$T_{f}/\log T_{f}=\Theta\left(\beta\xi_{u}+\xi_{w}/S\right)\,.$
@@
(D.15) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}T_{f}=\Theta\left((\beta\xi_{u}+\xi_{w}/S)\log\left(\beta\xi_{u}+\xi_{w}/S\right)\right)}\,.$
We can then estimate the integral by Eqs. D.12 and D.13, as well as Eq. D.15:
@@
$\int_{\mathbb{R}}\left|f(t)\right|\,\mathrm{d}t$ $=\int_{-T}^{T}|f(t)|\,\mathrm{d}t+\mathcal{O}\left(\xi^{2}_{q}S\int_{T^{1/s}}^{\infty}\exp\left(-\frac{s}{2e(\beta\xi_{u}+\xi_{w}/S)^{1/s}}u\right)\,\mathrm{d}u\right)$
(D.16) $=\mathcal{O}\left((\xi_{q}C_{1,u}+\xi^{2}_{q}\xi_{w})\log(T\max\{S,1\})+\xi^{2}_{q}\underbrace{S(\beta\xi_{u}+\xi_{w}/S)^{1/s}\exp\left(-\frac{sT^{1/s}}{2e(\beta\xi_{u}+\xi_{w}/S)^{1/s}}\right)}_{=:{\bf II}}\right)$
$=\mathcal{O}\left((\xi_{q}C_{1,u}+\xi^{2}_{q}\xi_{w})\log((\beta\xi_{u}+\xi_{w}/S)\max\{S,1\})\right),$
where $T=\Theta\left((\beta\xi_{u}+\xi_{w}/S)\log^{s}\left(\left(\beta\xi_{u}+\xi_{w}/S\right)\max\{S,1\}\right)\right)>T_{f}$ is chosen such that
${\bf II}$ $=S(\beta\xi_{u}+\xi_{w}/S)^{1/s}\exp\left(-\Theta(\log(\left(\beta\xi_{u}+\xi_{w}/S\right)\max\{S,1\}))\right)$
$=\frac{1}{(\beta\xi_{u}+\xi_{w}/S)^{\Theta(1)-1/s}}\frac{S}{\max\{S,1\}^{\Theta(1)}}\to 0\quad\text{as}\quad\beta\xi_{u}+\xi_{w}/S\to\infty\,.$
We have proved Eq. D.3.